The first of three numbers exceeds twice the second number by 4, while the third number is twice the first. If the sum of the three numbers is 54, find the number.
thanks to the person who will help me solve this question

Answer by ankor@dixie-net.com(22740)

:
The three numbers: x, y, z
Write an equation for each statement:
:
"The first of three numbers exceeds twice the second number by 4,"
x = 2y + 4
:
"the third number is twice the first."
z = 2x
:
"If the sum of the three numbers is 54,"
x + y + z = 54
:
find the number.
Try to get everything in terms of y:
z = 2x
Substitute (2y+4) for x (1st equation)
z = 2(2y+4)
z = (4y + 8)
:
In the 3rd equation substitute for x and z
x + y + z = 54
(2y+4) + y + (4y+8)
7y + 12 = 54
7y = 54 - 12
7y = 42
y = 42/7
y = 6
:
You should be able to find the values of x & z, using the above equation.
Check your solutions in the 3rd equation.
:
Any questions?